cos 2x + 2 cosx = 1 then evaluate sin2x (2 - cos2x).

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
cos2x + 2cosx = 1
=> 2cos2x – 1 + 2cosx = 1
=>  2cos2x + 2cosx = 2
=>  cos2x + cosx – 1 = 0
=> cosx = (-1 + √5)/2
From here we can calculate cos2x = 2cos2x – 1 and sin2x = √(1 – cos2 2x).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)